A man is 24 years older than his son.In two years,his age will be twicw the age of his son.the present age of the son is?
Answer by mananth(16949)

let son's age be x
father'a age = x+27
..
Age after 2 years
son = x+2
father = x+27+2
..
x+29 = 2(x+2)
x+29=2x+4
x=25 - son's age
Father's age = 25+27 = 52 years